# Demonstration of quadrature interferometric metrology of translation and tilt: QUIMETT

### v1

@@ Abstract
Future gravitational wave observation in space will demand the improvement on the sensitivity of the local sensor for the drag-free control. This paper presents the proposal, design, and demonstration of a new laser interferometric sensor named Quadrature Interferometric Metrology of Translation and Tilt (QUIMETT) for the drag-free local sensor. QUIMETT enables simultaneous measurements of both translational displacement and tilts of an reflective object with a single interferometer package. QUIMETT offers the characteristic feature where the sensitivity to tilt is independent of the interference condition while maintaining the ability to measure the translational displacement for a range greater than the laser wavelength. The tilt-sensing function has been demonstrated in a prototype experiment. The tilt sensitivity remained unchanged in different interference conditions and stayed at $10$ nrad/Hz1/2 at $0.1$ Hz.
  • October 31, 2024
@@ 1 Introduction
Gravitational waves were directly detected for the first time by Advanced Laser Interferometer Gravitational-wave Observatory (LIGO) in 2015 [1, 2]. Currently, the second half of fourth observing run, also known as O4b, by terrestrial gravitational wave antennas including LIGO, Virgo [3] are being performed, and KAGRA [4] will also be taking part in. In the mean time, space gravitational wave antennas including LISA [5], TianQin [6], Taiji [7], and DECIGO [8] have been in development or under consideration. The space gravitational wave antennas will probe gravitational waves oscillating at frequencies lower than a few Hz. Therefore, they are complementary to the terrestrial observatories. The space gravitational wave antennas are anticipated to promote the further research in the field of gravitational wave astronomy and astrophysics.
@@
The space gravitational wave antennas consist of inter-satellite laser interferometers with floating test masses aboard the satellites. The test masses serve as stable references for the gravitational wave observation. To isolate the test masses from noises associated with the random motion of satellite, the relative distance between the test mass and satellite has to be measured with local sensors and controlled by using the satellite thrusters. This scheme is known as the drag-free control. In the drag-free control, the sensing system often employs the capacitive sensors [9, 10] which offer a wide readout range and a decent displacement sensitivity of typically nm/Hz1/2 [11] with the hardware in a highly integrated form.
@@
For the space gravitational wave antennas seeking for a higher sensitivity, the improved sensitivity for the local sensors is necessary. For instance, DECIGO requires a displacement sensitivity of $10^{-12}$ m/Hz1/2 for the local sensors [12]. A promising sensor concept is the quadrature readout interferometer [13] which provides the displacement measurement of test mass for a range greater than the laser wavelength with a sensitivity better than the capacitive ones. In fact, there have been several research efforts developing the quadrature interferometers for precision measurement applications, such as Homodyne Quadrature Interferometer (HoQI) [14, 15], and modulation-assisted sensors [16, 17, 18, 19]. In addition, various solutions have been proposed to conduct multi-degrees-of-freedom measurement using interferometric sensors for broad applications  [20, 21, 22, 23, 24].
@@
Here, we propose a new sensor based on the HoQI scheme, Quadrature Interferometric Metrology of Translation and Tilt (QUIMETT). QUIMETT is a sensor with a simple configuration and no radio-frequency modulation is required, and is intended for use in space due to its resource-saving advantages with multi-degrees-of-freedom measurement in a single unit. QUIMETT adopts the linear combination of two tilt sensor outputs where the interference effect can be compensated to make the tilt sensing independent of the interference condition. This scheme makes QUIMETT capable of measuring the translation and two tilt degrees of freedom simultaneously. We have demonstrated the tilt sensing function and confirmed that the sensitivities do not vary as a function of the interference condition as expected. In this paper, we describe the working principle of the tilt measurement, design of a new optical system dedicated for the demonstration, and experimental verification.
@@
This paper is organized as follows. The principle of operation of the QUIMETT is presented in Section 2. The optical setup of the QUIMETT is given in Section 3. The experimental demonstration of the tilt sensing is presented in Section 4. After discussing the results in Section 5, we conclude this paper in Section 6.
@@
Figure 1: Conceptual setup of QUIMETT. The definition of the optical path lengths, $l$, and the Gouy phase shift along optical paths, $\eta$, are shown in Tab. 1. PD stands for photodetector, QPD for quadrant photodetector, NPBS for non-poarlizing beamsplitter, PBS for polarizing beamsplitter, HWP for half-wave plate, and QWP for quarter-wave plate.
@@ 2 Working principles
Figure 1 shows an illustration of conceptual setup for QUIMETT. We begin the analysis by computing the electric fields. Subsequently, the principles of translational displacement and tilt measurements are explained.
@@ 2.1 Electric fields in QUIMETT
where $A$ denotes polarization content i.e., $A\in(h,v)$ with $h$ and $v$ for horizontal and vertical polarizations, respectively. Horizontal polarization vibrates in perpendicular to this paper and vertical polarization vibrates in parallel to this paper. The electric field $E^{A}_{ij}$ represents the Hermite-Gauss contents for the $(i,j)$ mode. We hereafter denote the $6\times 6$ matrices and $6\times 1$ vectors in blackboard bold font while the $3\times 3$ matrices and $3\times 1$ vectors are written with superscripts hat and arrow, respectively.
We now derive the electric fields at the photo detectors namely, PD and QPDs 1 and 2 in Fig. 1, by letting the laser light propagate through the optics. The input field $\mathbb{E}_{\textrm{in}}$ denotes the one departing PBS1 for non-poarlizing beamsplitter (NPBS) and set as $\mathbb{E}_{\textrm{in}}=(0,0,0,\sqrt{P_{\textrm{in}}},0,0)^{\intercal}$ with $P_{\textrm{in}}$ the input power to the NPBS. Exerting the optical element matrices to the input field along the light paths, one can express the electric field vector at PD, $\mathbb{E}^{\textrm{PD}}$, as
@@
where $\mathbb{A}_{r}$, $\mathbb{B}_{A}$, $\mathbb{W}^{(2)}$ and $\mathbb{P}$ are the optical element matrices for non-polarizing reflection, polarizing beamsplitter (PBS), half-wave plate (HWP) and free-space propagation, respectively. The definition of those optical element matrices are given in A. $l$ and $\eta$ are the optical path lengths and the Gouy phase shift along optical paths, respectively. The detailed definitions of $l$ and $\eta$ are shown in Fig. 1 and Tab. 1. Here and hereafter, we assume that $l_{x}$ and $l_{y}$ are macroscopically the same and Gouy phase shifts in these path are the same, i.e. $\eta_{x_{\mathrm{in}}}=\eta_{y_{\mathrm{in}}}\equiv\eta_{\mathrm{ITF_{in}}}$ and $\eta_{x}=\eta_{y}\equiv\eta_{\mathrm{ITF}}$. $R_{\textrm{PBS2}}$ and $T_{\textrm{PBS2}}$ are reflectance for vertical polarization and transmittance for horizontal polarization of PBS2, respectively. The matrix $\mathbb{M}$ represents the reflection by the test mass with misalignment in both pitch and yaw directions by the normalized tilts $\Theta_{\textrm{yaw}}$ and $\Theta_{\textrm{pitch}}$ [25], respectively, as
@@
$\begin{split}\mathbb{E}^{\textrm{QPD}n}=&\mathbb{F}(n)\mathbb{W}^{\left(4\right)}\left(\frac{\pi}{4}\right)\mathbb{P}(l_{3},\eta_{3})\mathbb{A}_{\textrm{t}}\left(\frac{1}{2}\right)\mathbb{P}(l_{1},\eta_{1})\mathbb{W}^{\left(2\right)}\left(-\frac{\pi}{8}\right)\\ &\times\{\underbrace{\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw}},\Theta_{\textrm{pitch}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)}_{\textrm{test mass}}\\ &+\underbrace{\mathbb{B}_{h}(T_{\textrm{PBS2}})\mathbb{P}(l_{x},\eta_{\mathrm{ITF}})\mathbb{A}_{\textrm{r}}(1)\mathbb{P}(l_{x},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{h}\left(T_{\textrm{PBS2}}\right)}_{\textrm{reference mirror}}\}\mathbb{W}^{\left(2\right)}\left(\frac{\pi}{8}\right)\mathbb{P}(l_{1},\eta_{1_{\mathrm{in}}})\mathbb{A}_{\textrm{r}}\left(\frac{1}{2}\right)\mathbb{E}_{\textrm{in}},\end{split}$ (4)
where $\mathbb{A}_{\textrm{t}}$ is the optical element matrices for non-polarizing transmittance, $\mathbb{W}^{(4)}$ represents a quater-wave plate as defined in A, and we have introduced $\mathbb{F}$ for selecting the light path between those for QPDs 1 and 2. The definition is given as
$\mathbb{F}(1)=\mathbb{P}(l_{4},\eta_{4})\mathbb{B}_{h}(1),\quad\textrm{and}\quad\mathbb{F}(2)=\mathbb{P}(l_{5},\eta_{5})\mathbb{A}_{\textrm{r}}(1)\mathbb{B}_{v}(1).$ (5)
Here, we assume that the reflectance and transmittance of PBS3 are unity.
@@ 2.2 Translational displacement readout
$I_{\textrm{PD}}$ $=\frac{1-a\cos{2k(l_{x}-l_{y})}}{16}\left(R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}\right)P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (6)
$I_{\textrm{QPD1}}$ $=\frac{1-a\sin{2k(l_{x}-l_{y})}}{16}\left(R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}\right)P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (7)
$I_{\textrm{QPD2}}$ $=\frac{1+a\sin{2k(l_{x}-l_{y})}}{16}\left(R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}\right)P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (8)
@@
and retrieve the displacement by applying the arc-tangent operation. This highlights the advantage of quadrature readout interferometer where the range of translational displacement can be expanded to a range greater than the wavelength via an additional signal processing. QUIMETT maintains this ability since $I_{\textrm{QPD1}}$ ($I_{\textrm{QPD2}}$) can be measured with the sum signal of all segments in QPD1 (QPD2).
@@ 2.3 Tilt readout
The QPDs play a crucial role in deriving the information associated with the tilts. The tilt information for the yaw (pitch) direction can be readily obtained by taking the difference in light intensities between the left and right (upper and lower) segments. This results in tilt signal in the form of
@@
where $B$ is any tilt in pitch or yaw, and Eqs. (10) and (11) give the response to the tilt in the direction selected at $B$.
@@
Figure 2: Optical layout used in the experiment. A photograph in the upper right corner shows the actual setup. The actuated mirror at the bottom simulates the tilt of test mass while the one on the right of PBS adjusts the interference condition. The yellow texts and accompanying arrows illustrate the direction of motion for the actuator coupled to the mirrors.
@@ 3 Experimental setup
A convex lens is inserted in the optical system for two reasons. First, the lens keeps the beam size small enough throughout the system so that it does not require an additional train of lenses. The other reason is that it adjusts the Gouy phase evolution. As discussed in Sec. 2, the response to tilt deviations depends on the Gouy phase accumulated from the test mass to the locations of QPDs. If not carefully designed, the tilt signal would vanish in the worst case. We avoid such a situation by designing the beam waist locating in the vicinity of the test mass. In our experiment, the beam waist is measured to be 15 mm behind the test mass. The Gouy phase from the test mass to the QPDs is measured to be $106$ deg.
All the QPDs and PDs are connected to the transimpedance amplifiers (TIAs) via lead wires. The TIAs are mounted on a separate location. The voltage signals coming out of the TIAs are continuously recorded in time series. The measurement instrument is YOKOGAWA DL850E with 16bit modules.

### v2

# Demonstration of tilt sensing using a homodyne quadrature interferometric translational sensor

@@ Abstract
Future gravitational wave observation in space will demand improvement in the sensitivity of the local sensor for the drag-free control. This paper presents the proposal, design, and demonstration of a new laser interferometric sensor named Quadrature Interferometric Metrology of Translation and Tilt (QUIMETT) for the drag-free local sensor. QUIMETT enables simultaneous measurements of both translational displacement and tilts of a reflective object with a single interferometer package. QUIMETT offers a characteristic feature where the sensitivity to tilt is independent of the interference condition while maintaining the ability to measure the translational displacement for a range greater than the laser wavelength. The tilt-sensing function has been demonstrated in a prototype experiment. The tilt sensitivity remained unchanged in different interference conditions and stayed at $10$ nrad/Hz1/2 at $0.1$ Hz.
  • May 7, 2025
@@ 1 Introduction
Gravitational waves were directly detected for the first time by Advanced Laser Interferometer Gravitational-wave Observatory (LIGO) in 2015 [1, 2]. Currently, the second half of the fourth observing run, also known as O4b, by terrestrial gravitational wave antennas including LIGO, Virgo [3] is being performed, and KAGRA [4] will also be taking part in the observation run. In the meantime, space gravitational wave antennas including LISA [5], TianQin [6], Taiji [7], and DECIGO [8] have been in development or under consideration. The space gravitational wave antennas will probe gravitational waves oscillating at frequencies lower than a few Hz. Therefore, they are complementary to the terrestrial observatories. The space gravitational wave antennas are anticipated to promote the further research in the field of gravitational wave astronomy and astrophysics.
@@
The space gravitational wave antennas consist of inter-satellite laser interferometers with floating test masses aboard the satellites. The test masses serve as stable references for the gravitational wave observation. To isolate the test masses from noise associated with the random motion of the satellite, the relative distance between the test mass and satellite has to be measured with local sensors and controlled by using the satellite thrusters. This scheme is known as the drag-free control. In the drag-free control, the sensing system often employs the capacitive sensors [9, 10] which offer a wide readout range and a decent displacement sensitivity of typically nm/Hz1/2 [11] with the hardware in a highly integrated form.
@@
For the space gravitational wave antennas seeking for a higher sensitivity, the improved sensitivity for the local sensors is necessary. For instance, DECIGO requires a displacement sensitivity of $10^{-12}$ m/Hz1/2 for the local sensors [12]. A promising sensor concept is the quadrature readout interferometer [13] which provides the displacement measurement of the test mass for a range greater than the laser wavelength with a sensitivity better than the capacitive ones. In fact, there have been several research efforts developing the quadrature interferometers for precision measurement applications, such as Homodyne Quadrature Interferometer (HoQI) [14, 15], and modulation-assisted sensors [16, 17, 18, 19]. In addition, various solutions have been proposed to conduct multi-degrees-of-freedom measurement using interferometric sensors for broad applications  [20, 21, 22, 23, 24].
@@
Here, we propose a new sensor based on the HoQI scheme, Quadrature Interferometric Metrology of Translation and Tilt (QUIMETT). QUIMETT is a sensor with a simple configuration and no radio-frequency modulation is required, and is intended for use in space due to its resource-saving advantages with multi-degrees-of-freedom measurement in a single unit. QUIMETT adopts a linear combination of two tilt sensor outputs where the interference effect can be compensated to make the tilt sensing independent of the interference condition. This scheme makes QUIMETT capable of measuring the translation and two tilt degrees of freedom simultaneously. We have demonstrated the tilt sensing function and confirmed that the sensitivities do not vary as a function of the interference condition as expected. In this paper, we describe the working principle of the tilt measurement, design of a new optical system dedicated for the demonstration, and experimental verification.
@@
This paper is organized as follows. The principle of operation of the QUIMETT is presented in Section 2. The optical setup of QUIMETT is given in Section 3. The experimental demonstration of the tilt sensing is presented in Section 4. After discussing the results in Section 5, we conclude this paper in Section 6.
@@
Figure 1: Conceptual setup of QUIMETT. The definition of the optical path lengths, $l$, and the Gouy phase shift along optical paths, $\eta$, are shown in Tab. 1. PD stands for photodetector, QPD for quadrant photodetector, NPBS for non-polarizing beamsplitter, PBS for polarizing beamsplitter, HWP for half-wave plate, and QWP for quarter-wave plate.
@@ 2 Working principles
Figure 1 shows an illustration of the conceptual setup for QUIMETT. We begin the analysis by computing the electric fields. Subsequently, the principles of translational displacement and tilt measurements are explained.
@@ 2.1 Electric fields in QUIMETT
where $A$ denotes polarization content i.e., $A\in(h,v)$ with $h$ and $v$ for horizontal and vertical polarizations, respectively. Horizontal polarization is defined as vibrating within the $x$-$y$ plane, while vertical polarization is aligned along the $z$-axis. The electric field $E^{A}_{ij}$ represents the Hermite-Gauss contents for the $(i,j)$ mode. We hereafter denote the $6\times 6$ matrices and $6\times 1$ vectors in blackboard bold font while the $3\times 3$ matrices and $3\times 1$ vectors are written with superscripts hat and arrow, respectively.
We now derive the electric fields at the photodetectors namely, PD and QPDs 1 and 2 in Fig. 1, by letting the laser light propagate through the optics. The input field $\mathbb{E}_{\textrm{in}}$ denotes the one departing PBS1 for non-polarizing beamsplitter (NPBS) and is set as $\mathbb{E}_{\textrm{in}}=(0,0,0,\sqrt{P_{\textrm{in}}},0,0)^{\intercal}$ with $P_{\textrm{in}}$ the input power to the NPBS. Exerting the optical element matrices to the input field along the light paths, one can express the electric field vector at PD, $\mathbb{E}^{\textrm{PD}}$, as
@@
where $\mathbb{A}_{r}$, $\mathbb{B}_{A}$, $\mathbb{W}^{(2)}$ and $\mathbb{P}$ are the optical element matrices for non-polarizing reflection, polarizing beamsplitter (PBS), half-wave plate (HWP) and free-space propagation, respectively. The definitions of these optical element matrices are given in A. $l$ and $\eta$ are the optical path lengths and the Gouy phase shift along optical paths, respectively. The detailed definitions of $l$ and $\eta$ are shown in Fig. 1 and Tab. 1. Here and hereafter, we assume that $l_{x}$ and $l_{y}$ are macroscopically the same and Gouy phase shifts in these paths are the same, i.e. $\eta_{x_{\mathrm{in}}}=\eta_{y_{\mathrm{in}}}\equiv\eta_{\mathrm{ITF_{in}}}$ and $\eta_{x}=\eta_{y}\equiv\eta_{\mathrm{ITF}}$. $R_{\textrm{PBS2}}$ and $T_{\textrm{PBS2}}$ are reflectance for vertical polarization and transmittance for horizontal polarization of PBS2, respectively. The matrix $\mathbb{M}$ represents the reflection by the test mass with misalignment in both pitch and yaw directions by the normalized tilts $\Theta_{\textrm{yaw}}$ and $\Theta_{\textrm{pitch}}$ [25], respectively, as
@@
$\begin{split}\mathbb{E}^{\textrm{QPD}n}=&\mathbb{F}_{n}\mathbb{W}^{\left(4\right)}\left(\frac{\pi}{4}\right)\mathbb{P}(l_{3},\eta_{3})\mathbb{A}_{\textrm{t}}\left(\frac{1}{2}\right)\mathbb{P}(l_{1},\eta_{1})\mathbb{W}^{\left(2\right)}\left(-\frac{\pi}{8}\right)\\ &\times\{\underbrace{\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw}},\Theta_{\textrm{pitch}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)}_{\textrm{test mass}}\\ &+\underbrace{\mathbb{B}_{h}(T_{\textrm{PBS2}})\mathbb{P}(l_{x},\eta_{\mathrm{ITF}})\mathbb{A}_{\textrm{r}}(1)\mathbb{P}(l_{x},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{h}\left(T_{\textrm{PBS2}}\right)}_{\textrm{reference mirror}}\}\mathbb{W}^{\left(2\right)}\left(\frac{\pi}{8}\right)\mathbb{P}(l_{1},\eta_{1_{\mathrm{in}}})\mathbb{A}_{\textrm{r}}\left(\frac{1}{2}\right)\mathbb{E}_{\textrm{in}},\end{split}$ (4)
where $\mathbb{A}_{\textrm{t}}$ is the optical element matrices for non-polarizing transmittance, $\mathbb{W}^{(4)}$ represents a quarter-wave plate as defined in A, and we have introduced $\mathbb{F}_{n}\ (n=1,2)$ for selecting the light path between those for QPDs 1 and 2. The definition is given as
$\mathbb{F}_{1}=\mathbb{P}(l_{4},\eta_{4})\mathbb{B}_{h}(1),\quad\textrm{and}\quad\mathbb{F}_{2}=\mathbb{P}(l_{5},\eta_{5})\mathbb{A}_{\textrm{r}}(1)\mathbb{B}_{v}(1).$ (5)
Here, we assume that the reflectance and transmittance of PBS3 are unity. Note that proper control of the Gouy phase is crucial in our measurement approach, as it directly influences the response to tilt deviations.
@@ 2.2 Translational displacement readout
$I_{\textrm{PD}}$ $=\frac{1-a\cos{2k(l_{x}-l_{y})}}{8}\frac{R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}}{2}P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (6)
$I_{\textrm{QPD1}}$ $=\frac{1-a\sin{2k(l_{x}-l_{y})}}{8}\frac{R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}}{2}P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (7)
$I_{\textrm{QPD2}}$ $=\frac{1+a\sin{2k(l_{x}-l_{y})}}{8}\frac{R_{\textrm{PBS2}}^{2}+T_{\textrm{PBS2}}^{2}}{2}P_{\textrm{in}}+\mathcal{O}(\Theta^{2}),$ (8)
@@
and retrieve the displacement by applying the arc-tangent operation. This highlights the advantage of the quadrature readout interferometer where the range of translational displacement can be expanded to a range greater than the wavelength via additional signal processing. QUIMETT maintains this ability since $I_{\textrm{QPD1}}$ ($I_{\textrm{QPD2}}$) can be measured with the sum signal of all segments in QPD1 (QPD2).
@@ 2.3 Tilt readout
The QPDs play a crucial role in deriving the information associated with the tilts. The tilt signal for the yaw (pitch) direction can be readily obtained by taking the difference in light intensities between the left and right (upper and lower) segments. This results in tilt signals in the form of
@@
where $B$ is any tilt in pitch or yaw, and Eqs. (10) and (11) give the response to the tilt in the direction selected at $B$. In deriving Eqs. (10) and (11), we assume that the beam displacement from the center of the QPDs is small for both the reference and target beams. We also assume that the beam diameter is significantly smaller than the QPDs size, allowing us to neglect truncation effects. Additionally, we consider a setup without a lens between the target mirror and the QPDs, so the beam undergoes natural diffraction as it propagates, and its wavefront evolves without additional focusing effects.
@@
Figure 2: Optical layout used in the experiment. A photograph in the upper right corner shows the actual setup. The actuated mirror at the bottom simulates the tilt of the test mass while the one on the right of the PBS adjusts the interference condition. The yellow texts and accompanying arrows illustrate the direction of motion for the actuator coupled to the mirrors.
In our formulation, the dependence of light intensities on the distance from the target mirror to the QPD and beam width at the QPD is implicitly included through the Gouy phase rather than explicitly expressed in terms of the geometric factor. This approach provides a more general description, particularly when the system size is comparable to or larger than the Rayleigh range.
@@ 3 Experimental setup
A convex lens is inserted in the optical system for two reasons. First, the lens keeps the beam size small enough throughout the system so that it does not require an additional train of lenses. The other reason is that it adjusts the Gouy phase evolution. As discussed in Sec. 2, the response to tilt deviations depends on the Gouy phase accumulated from the test mass to the locations of QPDs. If not carefully designed, the tilt signal would vanish in the worst case. We avoid such a situation by designing the beam waist to be located in the vicinity of the test mass. In our experiment, the beam waist position and the waist diameter are measured to be 15 mm behind the test mass and $0.2$ mm, respectively. The Gouy phase from the test mass to the QPDs is measured to be $124$ deg. Note that the beam diameter on the QPDs is $0.7$ mm.
All the QPDs and PDs are connected to the transimpedance amplifiers (TIAs) via lead wires. The TIAs are mounted on a separate location. The voltage signals coming out of the TIAs are continuously recorded in a time series. The measurement instrument is YOKOGAWA DL850E with 16-bit modules.